ICM 1998, Berlin, Aug. 18–27

Abstracts of Plenary and Invited Lectures

Section:

12. Probability and Statistics

1991 MS Classification: 60K40,82D30 Sznitman, Alain-Sol, ETH Zurich, Switzerland Brownian motion and random obstacles

The investigation of Brownian motion and random obstacles exhibits a rich phenomenology and displays paradigms which appear in several other areas of random media. The lecture will try to give an overview of recent developments such as the pinning effect, or the confinement property. It will discuss some of the techniques which have emerged such as the method of enlargement of obstacles, or the Lyapunov exponents. It will also point out connections with other areas of random media.

ICM 1998 Berlin · Plenary and Invited Lectures